If y varies inversely as the square of x and y = 1 when x = 2, what is the value of y when x = -4?
Answer by nyc_function(2741)

y = k/x^2
We need to find k first.
Let y = 1, when x = 2 to find k.
1 = k/(2)^2
1 = k/4
1*4 = k
4 = k
We now need to find y when k = 4 and x = -4.
y = k/x^2
y = 4/(-4)^2
y = 4/16
y = 1/4
Done!